There are 2 friends. One's age is 4 years older than half his friend's age. Together their ages are 28. How old are they?
Answer by addingup(3677)

2 friends, we'll call the x and y
x= 1/2y+4 We'll use this value for x next:
x+y= 28 Substitute:
1/2y+4+y= 28 Multiply both sides times 2
y+8+2y= 56 Add y on left and subtract 8 on both sides
3y= 48 divide both sides by 3
y= 16 And the problem says that x is:
x= 1/2y+4; x= 1/2(16)+4; x= 8+4; x= 12
One friend is 16 and the other 12
Check:
16+12= 28 just like the problem says.